OpenStudy (anonymous):

f(x)= 3/x+6 g(x)=x^2-3 find(f/g)(x)

OpenStudy (anonymous):

\[f(x)=\frac{ 3 }{ x+6 }\] and \[g(x) =x^{2}-3\] ?????

OpenStudy (anonymous):

yes thats it

OpenStudy (jdoe0001):

\(\bf \left(\cfrac{f}{g}\right)x \implies \cfrac{f(x)}{g(x)}\\ \textit{keep in mind that } \cfrac{\frac{a}{b}}{\frac{c}{d}} \implies \cfrac{a}{b}\implies\cfrac{d}{c}\)

OpenStudy (anonymous):

i know how its set up i just dont know how to solve

OpenStudy (jdoe0001):

ohh ... darn.. .I have a typo

OpenStudy (jdoe0001):

\(\bf \left(\cfrac{f}{g}\right)x \implies \cfrac{f(x)}{g(x)}\\ \textit{keep in mind that } \cfrac{\quad\frac{a}{b}\quad }{\frac{c}{d}} \implies \cfrac{a}{b}\times\cfrac{d}{c}\)

OpenStudy (anonymous):

\[\frac{ f(x) }{g(x) } = \frac{ \frac{ 3 }{x + 6 } }{ x ^{2}-3 } = \frac{ 3 }{(x+6)(x ^{2}-3) }\]

OpenStudy (anonymous):

same thing jdoe0001

OpenStudy (anonymous):

so i should end up with \[\frac{ 3 }{ x ^{3}+6x ^{2}-3x-18 }\]

OpenStudy (anonymous):

it's ok